Circular No. 5929

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SCHWASSMANN-WACHMANN 1 M. C. Senay and D. Jewitt, University of Hawaii, report: “We have

detected CO J(2-1) emission at 230 GHz from P/Schwassmann-Wachmann 1 on 1993 Oct. 22, Nov. 11 and 12 UT. The observations were taken using the JCMT on Mauna Kea, and give a line strength of 0.08 ± 0.01 K km s^{−1}. A preliminary estimate of the CO production rate gives 2000 kg/s, which is sufficient to drive the dust activity observed in this comet. This observation provides the first direct evidence that activity in comets beyond the orbit of Jupiter can be powered by CO.”

GRO J0422+32 P. Zhao, P. Callanan, M. Garcia, and J. McClintock, Smithsonian

Astrophysical Observatory, write: “On Jan. 30.29 UT, the x-ray nova J0422+32 was fainter than at any time since its outburst in Aug. 1992 (IAUC 5580): R = 20.03 ± 0.11 and V = 20.67 ± 0.22, which is 8 mag fainter than its maximum brightness during outburst (IAUC 5588). These CCD photometric data were obtained with the F. L. Whipple Observatory’s 1.2-m telescope. The magnitudes on two preceding nights were: Jan. 21.15, R = 18.66 ± 0.03, V = 19.31 ± 0.06; 22.15, 19.09 ± 0.16, 19.83 ± 0.29. This new faint state provides the best opportunity to date of measuring the radial velocity variations of the secondary star.”

TIME ADJUSTMENT ON 1994 JUNE 30 Bulletin C7 of the International Earth Rotation Service announces

that a positive leap second will be introduced such that the sequence of UTC second markers will be: 1994 June 30^{d}23^{h}59^{m}59^{s}, 30^{d}23^{h}59^{m}60^{s}, July 1^{d}00^{h}00^{m}00^{s}. Beginning 1994 July 1, the difference UTC–TAI = −29^{s}.

NOVA CASSIOPEIAE 1993 Further photoelectric photometry (cf. IAUC 5920, 5928): Jan. 24.58

UT, V = 7.56, B–V = +0.42, U–B = −0.65 (B. N. Ashoka, Bangalore, India; comparison stars HR 9010, HR 8985); 25.58, 7.56, +0.41, −0.61 (Ashoka); 26.58, 7.70, +0.43, −0.62 (Ashoka); 28.59, 7.74, +0.44, −0.64 (Ashoka); 29.58, 7.95, +0.41, −0.61 (Ashoka); 29.80, 8.05, –, – (H. Mikuž, Ljubljana, Slovenia); 30.58, 8.08, +0.45, −0.65 (Ashoka); 30.75, 8.13, –, – (Mikuž); 31.58, 7.91, +0.44, −0.61 (Ashoka).

1994 February 4 Daniel W. E. Green